# Improving the Evaluation and Actionability of Explanation Methods for Multivariate Time Series Classification

### v1

@@ Abstract
Explanation for Multivariate Time Series Classification (MTSC) is an important topic that is under explored. There are very few quantitative evaluation methodologies and even fewer examples of actionable explanation, where the explanation methods are shown to objectively improve specific computational tasks on time series data. In this paper we focus on analyzing InterpretTime, a recent evaluation methodology for attribution methods applied to MTSC. We reproduce the original paper results, showcase some significant weaknesses of the methodology and propose ideas to improve both its accuracy and efficiency. Unlike related work, we go beyond evaluation and also showcase the actionability of the produced explainer ranking, by using the best attribution methods for the task of channel selection in MTSC. We find that perturbation-based methods such as SHAP and Feature Ablation work well across a set of datasets, classifiers and tasks and outperform gradient-based methods. We apply the best ranked explainers to channel selection for MTSC and show significant data size reduction and improved classifier accuracy.
@@ 1 Introduction
  • We quantitatively compare multiple explanation methods for MTSC, starting from an existing methodology named InterpretTime. We show that InterpretTime has some significant weaknesses with respect to the explainer ranking methodology, which fails for some datasets and has poor efficiency in obtaining the explainer evaluation.
  • We propose improvements to InterpretTime: we show how grouping consecutive time points channel-wise helps reduce running time and improve the performances of explainers.
@@ 2.3 Quantitative Evaluation of Attribution Methods for MTSC
Quantitative evaluation of explanation methods includes quantitative metrics such as robustness, stability, sparsity [9, 12] and perturbation-based frameworks such as AMEE [20, 21] and InterpretTime [29]. Robustness addresses the model’s capacity to deal with noise while sparsity measures how focused the saliency map is. Unfortunately, none of the above metrics give sufficient information on whether a particular saliency map is even better than a random one. AMEE and InterpretTime present another approach where noise is inserted into the input time series at different levels and under the guidance of the saliency map. The drop in the accuracy of the classifiers (due to mislabelling the corrupted samples) indicates the quality of the explanation. While InterpretTime works for MTSC, AMEE is designed for UTSC but also works with MTS data by concatenating the channels [24]. In this paper, we reproduce the results of the InterpretTime framework for MTSC and analyse this approach further.
@@ 3 Background
A multivariate time series can be presented as a $d\times L$ matrix $X$ where $d$ is the number of channels and $L$ is the length of the time series. Multivariate time series classification (MTSC) is the task of labeling unseen multivariate time series. The saliency map (or attribution) of $X$ is also a matrix $W$ of the same shape. The value $W_{j}^{i}$ at position $i,j$ ($j$-th item in $i$-th channel) indicates the importance of the corresponding input item ($X_{j}^{i}$) in the classification model. As the range of $W_{j}^{i}$ values can be different between methods, we used to normalize saliency maps in the range $[0,1]$.
@@ 3.4 Datasets
Synthetic data. The original dataset from [29] is composed of 6 different channels each having 500 time points and containing different sine waves. The discriminative features come from 2 randomly selected channels, in $100$ consecutive time points. In these positions, two other waves of higher frequency $f_{1}$ and $f_{2}$ replace the original ones. Each of the other channels has a $0.5$ probability of having a non-discriminative square wave. This binary classification task predicts whether or not $f_{1}+f_{2}$ is above a given threshold. In our implementation, two non-discriminative channels (number $6$ and $7$) are added; the test set is perfectly balanced having 500 positive and 500 negative instances, while the train set contains $7500$ samples with $3646$ positive ones.
@@ 4.2 Assessment based on Ground Truth
For the synthetic data we used an assessment based on the ground truth; we describe here its steps. The first one is to rescale the attribution matrices as they can have different magnitudes among different instances and different methods. To compare results with InterpretTime’s original results, we took for each attribution $W$, the $max(0,W_{j}^{i})$ (i.e., we consider only the positive values) and then normalize it in the range $[0,1]$.
The second step is to instantiate an explanation ground truth matrix $\mathcal{G}$ for each instance in the test set and compare the explanation against it. This is a matrix where all items are set to 00 except the ones corresponding to the discriminative elements in the time series, having values set to $1$. In other words, this is the ideal saliency map that can be achieved from any attribution method. The last step is to compute scores, namely PR-AUC and ROC-AUC [1] comparing each saliency map against the corresponding $\mathcal{G}$.
@@ 4.3 Improving InterpretTime
InterpretTime only uses values sampled from the $N(0,1)$ distribution to mask TS time points. We empirically verified that only relying on this distribution sometimes leads to a meaningless rank. This problem is due to the distribution shift from the train to the test set which in the original work was tackled by adding noise in the training procedure (more details in [29]). We instead believe this is a downside because requiring a specific training procedure limits the classifiers that can be applied, e.g., we cannot use pre-trained classifiers. We instead fixed this problem by using multiple masks. Other than $N(0,1)$, we also use Local and Global Gaussian and Local and Global mean as done in AMEE [20] plus the zeros distribution. Substituting the element $x_{t}^{c}$ (i.e., channel $c$, $t$-th time point of one element $x$ of the test set $X$), the replacing value $r_{t}^{c}$ is $r_{t}^{c}=0$ for the distribution of the zeros, while the others depend on the following means and standard deviations:
@@ 5.1 Validating InterpretTime Results
The first experiment was to validate the results presented in [29] using synthetic data. The InterpretTime methodology assesses the quality of saliency maps only by perturbing the time series, and the original study does not present a comparison to ground truth explanations for the synthetic data. We compared the explanation methods using the ground truth methodology illustrated in Section 4.2 (results are reported in Table 3).
Focusing on the XAI methods also tested in [29], we also found that SHAP is generally the best method (in our case for 3 out of 4 classifiers), KernelShap achieved all the time results very close to the random method and all gradient-based methods are in the middle between these two extremes. Differently from the original work we also found that SHAP has some of the poorest metrics using ConvTran, moreover GradientShap which is a competitive method in the original work, in our study is the second to last in both tested cases. In our experiments, we have also added FeatureAblation and FeaturePermutation, which can be seen as very simple and fast alternatives to SHAP. A final remark is that metrics are generally poor, suggesting that there is a weak overlap between the assessed saliency maps and the ground truth $\mathcal{G}$.
In our second experiment, we applied InterpretTime using only the $N(0,1)$ mask to compute the $AUC\tilde{S}_{top}$ score for each attribution, to directly compare with the original paper and with the metrics obtained using the ground truth. We were not able to reproduce the results from the original paper. Moreover, the ranking obtained using $AUC\tilde{S}_{top}$ is almost flat and very different than the one obtained using the metrics computed with ground truth. We show detailed results in the Appendix. In Section 5.2 we discuss what are the main sources of the problems pointed out by these two experiments (flat InterpretTime ranking and poor metrics for both cases) and the solutions we found for them.
@@ 5.2 Improved XAI Evaluation Methodology
Using only the $N(0,1)$ distribution, as the time series are progressively corrupted, all of the 4 classifiers predict the same class for every instance, specifically the negative one for synthetic data (more plots about classifiers’ behavior in the Appendix). Using this dataset, the $N(0,1)$ distribution, and the Local and Global Gaussian distributions are below the $15\%$ threshold, that is why they were excluded from scores computation, while the Global and Local mean and zeros distributions are above showing very close results. Table 4 reports the final metrics. We notice that for the best methods $AUC\tilde{S}_{top}$ is now remarkably higher than the corresponding random method, thus using the synthetic dataset our modified InterpretTime version leads to effective ranking, where metrics differences among different methods are now evident, and also have a better alignment with the ground truth results, especially for the ResNet classifier. In Section 5.3 another ranking problem due to the usage of $N(0,1)$ is shown. A final remark is that the $F1\mathcal{S}$ scores are most of the time proportional to $AUC\tilde{S}_{top}$: from now on we mostly focus on the $AUC\tilde{S}_{top}$ score in our analysis.
@@ 6 Conclusion
In this work we have validated the results of the recent InterpretTime methodology to compare different attribution methods for MTSC, using an assessment based on ground truth which was not considered in the original study. We found that for some dataset/classifier pairs the methodology cannot be run using the $N(0,1)$ distribution mask as it fails to provide a meaningful rank; we fixed this problem by introducing multiple masks and selecting only the ones producing a reasonable margin between the best and the random AUC score. We also improved the AUC scores of attribution methods, while reducing running time, by using chunking, i.e., grouping consecutive features in each channel. This is possible only for perturbation-based models as they can mask features together. Overall, SHAP is the best method with regard to accuracy. The simple FeatureAblation method is surprisingly not far from SHAP performance in terms of accuracy, but is much faster and overall better than other perturbation-based methods. We recommend using perturbation-based methods as the gradient-based methods are model-specific (deep learning) and cannot take into account the MTS locality as done with the chunking strategy. Finally, we run experiments focused on actionability, showing that the best attribution methods can correctly identify the most important channels in a dataset, using both synthetic and real-world datasets.
@@
More work is also required to better understand and address the mask-selection problem highlighted in the InterpretTime method. Selecting one single distribution for masking, regardless of whether it is $N(0,1)$ or something else, is to some extent arbitrary since is very unlikely to be completely uninformative. In this work we mitigated this problem by averaging up to 6 different masks but this point is still an open problem that requires more investigation.
@@ Acknowledgments
We would like to thank the anonymous reviewers for their detailed and constructive feedback. We want to thank all researchers working on time series and explainable AI who have made their data, code and results open source and have helped the reproducibility of research methods in this area. This work was funded by Science Foundation Ireland through the SFI Centre for Research Training in Machine Learning (18/CRT/6183) and the Insight Centre for Data Analytics (12/RC/2289_P2). For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@
Table 8: $AUC\tilde{S}_{top}$ and $F1\mathcal{S}$ for different attribution methods sorted by $AUC\tilde{S}_{top}$ , grouped by different classifiers. The whole pipeline (train the classifier, run XAI methods, run InterpretTime) was performed using our synthetic dataset implementation. All metrics, for each classifers, are very close to the random accuracy.
@@
Accuracy drop vs percentage of values substituted ($\tilde{N}$) curves for synthetic data using the original InterpretTime methodology (i.e. using values sampled from normal distribution as a mask). All explanation methods applied to all classifiers converge very fast to random accuracy ($0.5$ since this is a binary problem) because they classify all instances as negative ones.

### v2

@@ Abstract
Explanation for Multivariate Time Series Classification (MTSC) is an important topic that is under explored. There are very few quantitative evaluation methodologies and even fewer examples of actionable explanation, where the explanation methods are shown to objectively improve specific computational tasks on time series data. In this paper we focus on analyzing InterpretTime, a recent evaluation methodology for attribution methods applied to MTSC. We showcase some significant weaknesses of the original methodology and propose ideas to improve both its accuracy and efficiency. Unlike related work, we go beyond evaluation and also showcase the actionability of the produced explainer ranking, by using the best attribution methods for the task of channel selection in MTSC. We find that perturbation-based methods such as SHAP and Feature Ablation work well across a set of datasets, classifiers and tasks and outperform gradient-based methods. We apply the best ranked explainers to channel selection for MTSC and show significant data size reduction and improved classifier accuracy.
@@ 1 Introduction
  • We quantitatively compare multiple explanation methods for MTSC, starting from an existing methodology named InterpretTime. We show that InterpretTime has some significant weaknesses: it requires train data augmentation which limits the type of classifiers that can be used. Additionally, it has poor efficiency in obtaining the explainer evaluation.
  • We propose improvements to InterpretTime: We do not use data augmentation and instead use multiple masks. We also show how grouping consecutive time points channel-wise helps reduce running time and improve the performance of explainers.
@@ 2.3 Quantitative Evaluation of Attribution Methods for MTSC
Quantitative evaluation of explanation methods includes quantitative metrics such as robustness, stability, sparsity [9, 12] and perturbation-based frameworks such as AMEE [20, 21] and InterpretTime [29]. Robustness addresses the model’s capacity to deal with noise while sparsity measures how focused the saliency map is. Unfortunately, none of the above metrics give sufficient information on whether a particular saliency map is even better than a random one. AMEE and InterpretTime present another approach where noise is inserted into the input time series at different levels and under the guidance of the saliency map. The drop in the accuracy of the classifiers (due to mislabelling the corrupted samples) indicates the quality of the explanation. While InterpretTime works for MTSC, AMEE is designed for UTSC but also works with MTS data by concatenating the channels [24]. In this paper, we focus on the InterpretTime framework for MTSC and analyse this approach further.
@@ 3 Background
A multivariate time series can be presented as a $d\times L$ matrix $X$ where $d$ is the number of channels and $L$ is the length of the time series. Multivariate time series classification (MTSC) is the task of labeling unseen multivariate time series. The saliency map (or attribution) of $X$ is also a matrix $W$ of the same shape. The value $W_{j}^{i}$ at position $i,j$ ($j$-th item in $i$-th channel) indicates the importance of the corresponding input item ($X_{j}^{i}$) in the classification model. As the range of $W_{j}^{i}$ values can be different between methods, we normalize the saliency maps in the range $[0,1]$.
@@ 3.4 Datasets
Synthetic data. The original dataset from [29] is composed of 6 different channels each having 500 time points and containing different sine waves. The discriminative features come from 2 randomly selected channels, in $100$ consecutive time points. In these positions, two other waves of higher frequency $f_{1}$ and $f_{2}$ replace the original ones. Each of the other channels has a $0.5$ probability of having a non-discriminative square wave. This binary classification task predicts whether or not $f_{1}+f_{2}$ is above a given threshold. In our implementation we extend the public code to add two non-discriminative channels (number $6$ and $7$); the test set is perfectly balanced having 500 positive and 500 negative instances, while the train set contains $7500$ samples with $3646$ positive ones.
@@ 4.2 Assessment based on Ground Truth
For the synthetic data we used an assessment based on the ground truth; we describe here its steps. The first one is to rescale the attribution matrices as they can have different magnitudes among different instances and different methods. To compare results with InterpretTime’s original results, we took for each attribution $W$, the $max(0,W_{j}^{i})$ (i.e., we consider only the positive values) and then normalize it in the range $[0,1]$. The second step is to instantiate an explanation ground truth matrix $\mathcal{G}$ for each instance in the test set and compare the explanation against it. This is a matrix where all items are set to 00 except the ones corresponding to the discriminative elements in the time series, having values set to $1$. In other words, this is the ideal saliency map that can be achieved from any attribution method. The last step is to compute scores, namely PR-AUC and ROC-AUC [1] comparing each saliency map against the corresponding $\mathcal{G}$.
@@ 4.3 Improving InterpretTime
InterpretTime only uses values sampled from the $N(0,1)$ distribution to mask TS time points. We empirically verified that only relying on this distribution sometimes leads to a flat rank. This problem is due to the distribution shift from the train to the test set which in the original work was tackled by adding noise in the training procedure (more details in [29]). We believe this is a downside because requiring a specific training procedure limits the classifiers that can be applied, e.g., we cannot use pre-trained classifiers and it reduces the accuracy of non deep learning classifiers. We instead do not augment the training data as done in [29] and fixed this problem by using multiple masks. Besides $N(0,1)$, we use Local and Global Gaussian and Local and Global mean as done in AMEE [20] and the zeros distribution. Substituting the element $x_{t}^{c}$ (i.e., channel $c$, $t$-th time point of one element $x$ of the test set $X$), the replacing value $r_{t}^{c}$ is $r_{t}^{c}=0$ for the zeroes distribution, while the others depend on the following means and standard deviations:
@@ 5.1 Validating InterpretTime Results
The first experiment was to compare to the results presented in [29] using synthetic data. The InterpretTime methodology assesses the quality of saliency maps only by perturbing the time series, and the original study does not present a comparison to ground truth explanations for the synthetic data. We compared the explanation methods using the ground truth methodology illustrated in Section 4.2 (results are reported in Table 3).
Focusing on the XAI methods also tested in [29], we also found that SHAP is generally the best method (in our case for 3 out of 4 classifiers), KernelShap achieved all the time results very close to the random method and all gradient-based methods are in the middle between these two extremes. Different from the original work, we also found that SHAP has some of the poorest metrics using ConvTran, moreover GradientShap which is a competitive method in the original work, in our study is the second to last in both tested cases. In our experiments, we have also added FeatureAblation and FeaturePermutation, which can be seen as very simple and fast alternatives to SHAP. A final remark is that metrics are generally poor, suggesting that there is a weak overlap between the assessed saliency maps and the ground truth $\mathcal{G}$.
In our second experiment, we applied InterpretTime using only the $N(0,1)$ mask to compute the $AUC\tilde{S}_{top}$ score for each attribution, to directly compare with the original paper and with the metrics obtained using the ground truth. Under the setup we tested, namely no train data augmentation, we were not able to reproduce the results from the original paper. Moreover, the ranking obtained using $AUC\tilde{S}_{top}$ is almost flat and very different than the one obtained using the metrics computed with ground truth. We show detailed results in the Appendix. In Section 5.2 we discuss what are the main sources of the problems pointed out by these two experiments (flat InterpretTime ranking and poor metrics for both cases) and the solutions we found for them.
@@ 5.2 Improved XAI Evaluation Methodology
Using only the $N(0,1)$ distribution, as the time series are progressively corrupted, all of the 4 classifiers predict the same class for every instance, specifically the negative one for synthetic data (see Appendix). Using this dataset, the $N(0,1)$ distribution, and the Local and Global Gaussian distributions are below the $15\%$ threshold, that is why they were excluded from scores computation, while the Global and Local mean and zeros distributions are above, showing very close results. Table 4 reports the final metrics. We notice that for the best methods $AUC\tilde{S}_{top}$ is now remarkably higher than the corresponding random method, thus using the synthetic dataset our modified InterpretTime version leads to effective ranking, where metrics differences among different methods are now evident, and also have a better alignment with the ground truth results, especially for the ResNet classifier. In Section 5.3 another ranking problem due to the usage of $N(0,1)$ is shown. A final remark is that the $F1\mathcal{S}$ scores are most of the time proportional to $AUC\tilde{S}_{top}$: from now on we mostly focus on the $AUC\tilde{S}_{top}$ score in our analysis.
@@ 6 Conclusion
In this work we have investigated the recent InterpretTime methodology to compare different attribution methods for MTSC, using an assessment based on ground truth which was not considered in the original study. We found that, if we remove the train data augmentation used in the original study, for some dataset/classifier pairs the methodology cannot be run using only the $N(0,1)$ distribution mask as it provides a flat rank; we mitigated this problem by introducing multiple masks and selecting only the ones producing a reasonable margin between the best and the random AUC score. We also improved the AUC scores of attribution methods, while reducing running time, by using chunking, i.e., grouping consecutive features in each channel. This is possible only for perturbation-based models as they can mask features together. Overall, SHAP is the best method with regard to accuracy. The simple FeatureAblation method is surprisingly not far from SHAP accuracy, but is much faster and overall better than other perturbation-based methods. We recommend using perturbation-based methods as the gradient-based methods are model-specific (deep learning) and cannot take into account the MTS locality as done with the chunking strategy. Finally, we run experiments focused on actionability, showing that the best attribution methods can correctly identify the most important channels in a dataset, using both synthetic and real-world datasets.
@@
More work is also required to better understand and address the mask-selection problem highlighted in the InterpretTime method. Selecting one single distribution for masking is not enough, because different masks can lead to different ranking. In this work we mitigated this problem by averaging up to 6 different masks but this point is still an open problem that requires more investigation.
@@ Acknowledgments
We thank the anonymous reviewers for their constructive feedback. We are grateful to Jiawen Wei and Gianmarco Mengaldo for detailed discussions on the original InterpreTime methodology. We thank all researchers working on time series and explainable AI who have made their data, code and results open source to help the reproducibility of research methods in this area. This work was funded by Science Foundation Ireland through the SFI Centre for Research Training in Machine Learning (18/CRT/6183) and the Insight Centre for Data Analytics (12/RC/2289_P2). For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@
Table 8: $AUC\tilde{S}_{top}$ and $F1\mathcal{S}$ for different attribution methods sorted by $AUC\tilde{S}_{top}$ , grouped by different classifiers. The whole pipeline (train the classifier, run XAI methods, run InterpretTime with no data augmentation) was performed using our synthetic dataset implementation. All metrics, for each classifers, are very close to the random accuracy.
@@
Accuracy drop vs percentage of values substituted ($\tilde{N}$) curves for synthetic data using the original InterpretTime methodology (i.e. using values sampled from normal distribution as a mask, but no train data augmentation). All explanation methods applied to all classifiers converge very fast to random accuracy ($0.5$ since this is a binary problem) because they classify all instances as negative ones.